Some students planned for a get-together. The budget for food was $500. Five of the students failed to come because of the distance, therfore the cost of food for each member increased by $5. How many students attended the get-together?
Answer by checkley77(12569)

500/X=500/(X-5)-5
500/X=(500-5[X-5])/(X-5)
500/X=(500-5X+25]/(X-5)
X(500-5X+25)=500(X-5)
500X-5X^2+25X=500X-2500
-5X^2+25X+2500=0
-5(X^2-5X-500)=0
5(X-25)(X+20)=0
X-25=0
X=25 ANSWER.
PROOF:
500/25=500/(25-5)-5
20=500/20-5
20=25-5
20=20